# Machine Learning with Multitype Protected Attributes: Intersectional Fairness through Regularisation

### v1

@@ 2.2 Fairness gerrymandering
It is also important to note that fairness violations may arise even when protected attributes are excluded from the model inputs. This is due to proxy discrimination, where other variables that are strongly associated with protected attributes encode sensitive information. A well-known example is redlining, where geographic location serves as a proxy for ethnicity, leading to discriminatory outcomes (Prince and Schwarcz, 2019; Squires, 2003). In such cases, excluding protected attributes (e.g. ethnicity) does not eliminate the risk of unfairness, as proxy variables (e.g. geographic location) can still introduce disparities across different ethnicity groups. The risk of proxy discrimination is exacerbated by the power of modern ML algorithms. Promoting subgroup fairness helps mitigate this issue by controlling for dependencies between model predictions and the joint distribution of protected attributes, including those indirectly encoded through proxies.
@@ 2.3 Unfairness mitigation strategies
Approaches to mitigating unfairness in machine learning fall into three broad categories: pre-processing, in-processing, and post-processing methods (Mehrabi et al., 2021; Caton and Haas, 2024; Pessach and Shmueli, 2022; Hort et al., 2024; Rabonato and Berton, 2024). Pre-processing methods modify the input data to reduce bias before training, using techniques such as fair representation learning (Zemel et al., 2013; Creager et al., 2019; Guo et al., 2022) or re-sampling strategies (Calders et al., 2009; Chakraborty et al., 2020). Post-processing methods adjust model outputs after training, including decision threshold adjustments and distribution transformations (Kamiran et al., 2012; Pope and Sydnor, 2011; Chzhen et al., 2019; Lindholm et al., 2022; Zeng et al., 2024).
@@ 2.4 Current approaches and limitations
Another limitation is the inability to capture nonlinear dependencies. Approaches such as absolute correlation (Beutel et al., 2019), and correlation between predictions and input attributes closely related to the protected attributes (Zhao et al., 2022), rely on Pearson’s correlation or similar linear metrics. These methods do not ensure full independence between predictions and protected attributes. Moreover, correlation can only capture pairwise associations and cannot capture more complex dependencies or interactions (Johnson et al., 2002; Makowski et al., 2020), making it ineffective in real world settings where relationships are often nonlinear and multi-dimensional. It is also not suitable for scenarios involving multiple protected attributes, where fairness must be ensured across combinations of variables.
@@ 4.2.3 Distribution-based metrics
where $\pi_{s}$ is the proportion of samples in each subgroup, and the KL-divergence between two probability distributions $P$ and $Q$ is defined as $D_{KL}(P||Q)=\sum\limits_{x}P(x)\log\frac{P(x)}{Q(x)}$, which quantifies how much information is lost when $Q$ is used to approximate $P$ (Kullback and Leibler, 1951). For continuous protected attributes, the presence of infinitely many values leads to an uncountable number of conditional distributions, making direct computation of $D_{JS}$ challenging. As a practical remedy, we discretise the continuous attributes into intervals, enabling estimation of $D_{JS}$ over a finite set of subgroups. Importantly, since discretisation is a deterministic function of the original variables, independence between $\hat{\mathbf{Y}}$ and the original (continuous or joint) protected attributes implies independence from their discretised version. As a result, the JS-divergence computed from the discretised subgroups will also be zero in case of independence. A smaller $D_{JS}$ represents a more aligned prediction distributions across subgroups, indicating a fairer model.
@@ 4.3 Model calibration
First, we tune the hyperparameters of the neural network on the training set without regularisation ($\lambda=0$) to obtain a baseline model. This is done using 5-fold cross-validation with Gaussian-process Bayesian optimisation and early stopping within each fold. The result is a single set of baseline hyperparameters, which are then fixed and used in all subsequent experiments. Further details are provided in Appendix D.
@@ 4.3.1 Selecting fairness metrics for $\lambda$ calibration
By directly quantifying distributional similarity, JS-divergence (see Section 4.2.3) offers an intuitive and interpretable measure for assessing how closely subgroup predictions align with the overall distribution. In addition, excessively large values of $\lambda$ may distort the output distributions across protected subgroups due to numerical instability. Importantly, such distortions are not always evident from the size of the regulariser term itself: even when the regulariser appears small, the resulting subgroup distributions may still become dispersed. JS-divergence thus also serves as a valuable tool for detecting discrepancies caused by over-regularisation. Accordingly, we advocate its use as the primary criterion for calibrating regularisation strength.
@@ 4.3.2 Procedure for tuning regularisation strength
The choice of $\lambda$ is fully exogenous and has to be made by the modeller. Here we explain and illustrate how to make such a choice, but the appropriate level of regularisation ultimately depends on the broader fairness objectives of the application, such as ethical standards or regulatory requirements. The calibration range for $\lambda$ should reflect the importance of fairness in the specific context. Moreover, because (joint) distance covariance is not scale-invariant under different datasets even when input features are standardised, the range of candidate $\lambda$ values will need to be selected so as to ensure that the regularisation term meaningfully contributes to the loss function. One possible starting point is to select $\lambda$ values with magnitudes comparable to the ratio of the baseline model’s validation loss to the regulariser (JdCov or CCdCov) computed using the predictions of the baseline model. Specifically, this ratio is given by
@@ 5.1.1 Background
As introduced earlier, the Correctional Offender Management Profiling for Alternative Sanctions (COMPAS) tool is an assessment instrument developed by Northpointe to estimate a defendant’s likelihood of recidivism based on responses to a survey (Corbett-Davies et al., 2016). Although Northpointe asserts that COMPAS is race-neutral (Bornstein, 2021), ProPublica’s analysis revealed that African-American defendants were significantly more likely to be classified as having a high likelihood of recidivism compared to Caucasian defendants. In particular, African-American males had the highest likelihood of being incorrectly classified as high risk, raising serious concerns about potential racial bias in the COMPAS risk assessment (Angwin et al., 2016).
@@ 5.1.2 Data description
We next consider interactions between protected attributes, as they may reveal hidden sources of bias not apparent when attributes are examined individually. Figure 5.1(d) shows the average recidivism rates across subgroups defined by the combination of gender, race, and age. For the purpose of this exploratory analysis, age is categorised into three groups based on the 33rd and 67th percentiles of the overall age distribution: 0–27, 28–37, and 38 or above. Notably, African-American males under 27 years of age exhibit the highest average recidivism rates. Again, this pattern may be reflected in the model’s predictions when no fairness treatments are applied.
Figure 5.1: Average recidivism rate by protected attributes
@@ 5.2.4 Results and analysis
In insurance pricing, it is essential for the premium to be unbiased at portfolio level, such that the predicted premium should not over- or under-estimate the total claim cost on average (Lindholm and Wüthrich, 2024). However, it is likely for our model to be biased, given we have applied early stopping and regularisation to prevent overfitting and fairness intervention (Wüthrich, 2020). In that case, we suggest applying a bias correction method such as the simple post-processing method introduced by Lindholm and Wüthrich (2024) to ensure the balance property of our predicted premium, such that the sum of predicted premiums should be the same as the sum of the total claim cost, weighted by exposure. In theory, the balance property guarantees an unbiased premium for any future portfolio whose risk profile, that is, the distribution of rating factors, matches that of the training data.
@@ Acknowledgments
All results in this paper are reproducible using Python, with code and data available on Zenodo (link to be updated upon submission) and GitHub (https://github.com/agi-lab/dCov-fairness). The code was run in an Azure ML notebook on a Standard_E16s_v3 virtual machine with 16 cores, which supports the full pipeline but was used specifically for the hyperparameter tuning phase. The Azure notebook is available on Zenodo, while a Quarto-adapted version was used for the final runs and is provided on GitHub. For the COMPAS dataset, hyperparameter tuning took approximately 4 hours. For the pg15training dataset, tuning was stopped after 24 hours, as continuing would have required considerably more time and computational resources. This early stopping is not expected to materially affect the results, as the large dataset size supports good generalisation performance. The remaining parts of the pipeline (training and evaluation using the selected hyperparameters) were run in Quarto locally, taking approximately 5 hours for COMPAS and 7 hours for pg15training.
@@ Appendix A Details on distance covariance
where $\langle\mathbf{a},\mathbf{b}\rangle=\sum_{i=1}^{m}a_{i}b_{i}$ is the dot product of two vectors $\mathbf{a},\mathbf{b}\in\mathbb{R}^{m}$. Let $\mathbf{Y}$ be another random vector of dimension $n$. We say $\mathbf{X}$ and $\mathbf{Y}$ are independent if and only if their characteristic function satisfy
@@
which follows directly from the definition of independence (Billingsley, 2017). This motivates the development of a measure of distance between $\phi_{\mathbf{X}}(t)\phi_{\mathbf{Y}}(s)$ and $\phi_{\mathbf{X},\mathbf{Y}}(t,s)$. Let $f$ be a function with domain $\Omega$ and $w(\cdot)$ be a nonnegative weight function. The weighted $L_{2}$ norm of $f$ is defined as
@@
Using (A.3), Székely et al. (2007) developed the distance covariance which is the weighted $L_{2}$ norm of the difference between $\phi_{X,Y}(t,s)$ and $\phi_{X}(t)\phi_{Y}(s)$:
@@
The unbiased estimator of the distance covariance, computed from a sample of $\mathbf{X}$ and $\mathbf{Y}$, is defined as (Székely and Rizzo, 2014):
@@
with $|\cdot|$ denoting the Euclidean norm. The $\mathcal{U}$-centred matrix is a bias-corrected form of the pairwise distance matrix. It is specifically designed so that the resulting estimator of distance covariance is unbiased under independence, meaning its expectation is zero when the variables are independent (Székely and Rizzo, 2023). Furthermore, following the proof of Proposition 1 in (Székely and Rizzo, 2014), we can expand Equation (A.6) into an equivalent form:
@@ B.1 Mutual independence
Let $\mathbf{X}_{1},\ldots,\mathbf{X}_{d}$ be random vectors with dimensions $p_{1},\ldots,p_{d}$, taking values in $\mathbb{R}^{p_{1}},\ldots,\mathbb{R}^{p_{d}}$. Note that this also covers one-hot–encoded categorical variables, which can be treated as a multinoulli distribution (Bishop and Nasrabadi, 2006; Murphy, 2012).
@@
$p(\mathbf{X}_{i}|\boldsymbol{\mu})=\prod^{K}_{k=1}\mu_{k}^{x_{i,k}}.$ (B.1)
@@ B.2 Joint distance covariance
$\begin{aligned} &dCov^{2}(\mathbf{X}_{\tau_{1}},...,\mathbf{X}_{\tau_{k}})\\ &\qquad=\int_{\mathbb{R}^{p_{0}}}\left|\mathbb{E}\left[\prod^{k}_{i=1}\left(\phi_{\mathbf{X}_{\tau_{i}}}(\mathbf{x}_{\tau_{i}})-\exp\left\{\sqrt{-1}\langle\mathbf{x}_{\tau_{i}},\mathbf{X}_{\tau_{i}}\rangle\right\}\right)\right]\right|^{2}dw,\end{aligned}$ (B.4)
with $dw=\prod^{k}_{i=1}\left(c_{p_{i}}dx_{\tau_{i}}\right)$ and $\mathcal{I}_{k}=\left\{(\tau_{1},\ldots,\tau_{k})\in\{1,\ldots,d\}^{k}:\tau_{1}<\tau_{2}<\cdots<\tau_{d}\right\}$ being the set of index tuples of size $d$. Intuitively, it is the summation of the pairwise associations ($dCov^{2}(\cdot,\cdot,\cdots,\cdot)$) and all the higher-order associations ($dCov^{2}(\cdots)$) between the random vectors, and it characterises mutual independence if and only if this quantity is 0. To estimate this quantity with the sample, Chakraborty and Zhang (2019) defined the biased-corrected estimator of the joint distance covariance as:
@@
with $\mathcal{I}\subseteq\{1,\ldots,d\}$ and $|\mathcal{I}|\geq 3$. In terms of model predictions and $d\geq 2$ protected attributes, we have:
@@
where the term $\sum_{1\leq i,j\leq d}\widetilde{dCov}^{2}(\mathbf{s}_{i},\mathbf{s}_{j})+\zeta(\hat{\mathbf{y}},\mathbf{s}_{1},...,\mathbf{s}_{d})$ captures the association within the protected attributes.
@@
For notation simplicity, denote ${\mathbf{s}}=(\mathbf{s}_{1},\ldots,\mathbf{s}_{d})$ and ${\mathbf{s}}_{k}=(\mathbf{s}_{k,1},\ldots,\mathbf{s}_{k,d})$ be the $k$-th sample of $\mathbf{s}$}. By Equation (A.8), we have:
$\widetilde{dCov}^{2}(\hat{\mathbf{y}},{\mathbf{s}})=\frac{1}{n(n-3)}\Biggl{(}\sum_{k=1}^{n}\sum_{l=1}^{n}|\hat{\mathbf{y}}_{k}-\hat{\mathbf{y}}_{l}||{\mathbf{s}}_{k}-{\mathbf{s}}_{l}|$ (C.1)
@@
$\qquad-\biggl{.}\frac{2}{n-2}\sum_{k=1}^{n}\biggl{(}\sum^{n}_{l=1}|\hat{\mathbf{y}}_{k}-\hat{\mathbf{y}}_{l}|\biggr{)}\biggl{(}\sum^{n}_{l=1}|{\mathbf{s}}_{k}-{\mathbf{s}}_{l}|\biggr{)}\Biggr{)}.$
@@
The term $\eta(\hat{\mathbf{y}},\mathbf{s}_{1},\ldots,\mathbf{s}_{d})$ captures the dependence that only shows up when the protected features are considered together. It is the residual link between the prediction $\hat{\mathbf{Y}}$ and the joint vector $\mathbf{S}=(\mathbf{S}_{1},\ldots,\mathbf{S}_{d})$ after removing the sum of the marginal dependencies with each $S_{i}$. If $\eta$ is large, the prediction is reacting to intersections of attributes even if each attribute on its own looks nearly independent. If $\hat{\mathbf{Y}}$ is independent of the whole vector $\mathbf{S}$, then $\eta=0$, and $\hat{\mathbf{Y}}$ is also independent of every subset of protected features.
Appendix D Details on model architecture

### v2

@@ 2.2 Fairness gerrymandering
It is also important to note that fairness violations may arise even when protected attributes are excluded from the model inputs. This is due to proxy discrimination, where other variables that are strongly associated with protected attributes encode sensitive information. A well-known example is redlining, where geographic location serves as a proxy for ethnicity, leading to discriminatory outcomes (Prince and Schwarcz, 2020; Squires, 2003). In such cases, excluding protected attributes (e.g., ethnicity) does not eliminate the risk of unfairness, as proxy variables (e.g., geographic location) can still introduce disparities across different ethnicity groups. The risk of proxy discrimination is exacerbated by the power of modern ML algorithms. Promoting subgroup fairness helps mitigate this issue by controlling for dependencies between model predictions and the joint distribution of protected attributes, including those indirectly encoded through proxies.
@@ 2.3 Unfairness mitigation strategies
Approaches to mitigating unfairness in machine learning fall into three broad categories: pre-processing, in-processing, and post-processing methods (Mehrabi et al., 2021; Caton and Haas, 2024; Pessach and Shmueli, 2022; Hort et al., 2024; Rabonato and Berton, 2025). Pre-processing methods modify the input data to reduce bias before training, using techniques such as fair representation learning (Zemel et al., 2013; Creager et al., 2019; Guo et al., 2022) or re-sampling strategies (Calders et al., 2009; Chakraborty et al., 2020). Post-processing methods adjust model outputs after training, including decision threshold adjustments and distribution transformations (Kamiran et al., 2012; Pope and Sydnor, 2011; Chzhen et al., 2019; Lindholm et al., 2022; Zeng et al., 2024).
@@ 2.4 Current approaches and limitations
Another limitation is the inability to capture nonlinear dependencies. Approaches such as absolute correlation (Beutel et al., 2019), and correlation between predictions and input attributes closely related to the protected attributes (Zhao et al., 2022), rely on Pearson’s correlation or similar linear metrics. These methods do not ensure full independence between predictions and protected attributes. Moreover, correlation can only capture pairwise associations and cannot capture more complex dependencies or interactions (Johnson and Wichern, 2002; Makowski et al., 2020), making it ineffective in real world settings where relationships are often nonlinear and multi-dimensional. It is also not suitable for scenarios involving multiple protected attributes, where fairness must be ensured across combinations of variables.
@@ 4.2.3 Distribution-based metrics
where $\pi_{s}$ is the proportion of samples in each subgroup, and the KL-divergence between two probability distributions $P$ and $Q$ is defined as $D_{KL}(P||Q)=\sum\limits_{x}P(x)\log\frac{P(x)}{Q(x)}$, which quantifies how much information is lost when $Q$ is used to approximate $P$ (Kullback and Leibler, 1951). For continuous protected attributes, the presence of infinitely many values leads to an uncountable number of conditional distributions, making direct computation of $D_{JS}$ challenging. As a practical remedy, we discretise the continuous attributes into intervals, enabling estimation of $D_{JS}$ over a finite set of subgroups. Importantly, since discretisation is a deterministic function of the original variables, independence between $\hat{\mathbf{Y}}$ and the original (continuous or joint) protected attributes implies independence from their discretised version. As a result, the JS-divergence computed from the discretised subgroups will also be zero in case of independence. A smaller $D_{JS}$ represents more aligned prediction distributions across subgroups, indicating a fairer model.
@@ 4.3 Model calibration
First, we tune the hyperparameters of the neural network on the training set without regularisation ($\lambda=0$) to obtain a baseline model. Here, hyperparameters refer to model-architecture and training settings (e.g., number of layers/neurons per layer, dropout, learning rate, batch size). This is done using 5-fold cross-validation with Gaussian-process Bayesian optimisation and early stopping within each fold. The result is a single set of baseline hyperparameters, which are then fixed and used in all subsequent experiments. Further details are provided in Appendix D.
@@ 4.3.1 Selecting fairness metrics for $\lambda$ calibration
By directly quantifying distributional similarity, JS-divergence (see Section 4.2.3) offers an intuitive and interpretable measure for assessing how closely subgroup predictions align with the overall distribution. In addition, excessively large values of $\lambda$ may distort the output distributions across protected subgroups when the fairness term dominates optimisation. Importantly, such distortions are not always evident from the size of the regulariser term itself: even when the regulariser appears small after training (e.g., after early stopping), the resulting subgroup distributions may still become dispersed. JS-divergence thus also serves as a valuable tool for detecting discrepancies caused by over-regularisation. Accordingly, we advocate its use as the primary criterion for calibrating regularisation strength.
@@ 4.3.2 Procedure for tuning regularisation strength
The choice of $\lambda$ is fully exogenous and has to be made by the modeller. Here we explain and illustrate how to make such a choice, but the appropriate level of regularisation ultimately depends on the broader fairness objectives of the application, such as ethical standards or regulatory requirements. The calibration range for $\lambda$ should reflect the importance of fairness in the specific context. Moreover, because (joint) distance covariance is not scale-invariant under different datasets even when input features are standardised, the range of candidate $\lambda$ values will need to be selected so as to ensure that the regularisation term meaningfully contributes to the loss function. One possible starting point is to select $\lambda$ values with magnitudes comparable to the ratio of the baseline ($\lambda=0$) model’s validation loss to the regulariser (JdCov or CCdCov) computed using the predictions of the baseline model. Specifically, this ratio is given by
@@ 5.1.1 Background
As introduced earlier, the Correctional Offender Management Profiling for Alternative Sanctions (COMPAS) tool is an assessment instrument developed by Northpointe to estimate a defendant’s likelihood of recidivism based on responses to a survey (Corbett-Davies et al., 2016). Although Northpointe asserts that COMPAS is race-neutral (Bornstein, 2021), ProPublica’s analysis revealed that African-American defendants were significantly more likely to be classified as having a high likelihood of recidivism than Caucasian defendants. In particular, African-American males had the highest likelihood of being incorrectly classified as high risk, raising serious concerns about potential racial bias in the COMPAS risk assessment (Angwin et al., 2016).
Additionally, recent fairness concerns in the legal system have arisen over the use of algorithmic recidivism assessment tools in guiding judgements. Written warnings now explicitly acknowledge that COMPAS may “disproportionately classify minority offenders as having a higher risk of recidivism” (Cox, 2017). These concerns motivate us to mitigate unfairness in the predicted recidivism rates, while still maintaining adequate accuracy to ensure the predictions remain informative for guiding legal judgements.
@@ 5.1.2 Data description
We next consider interactions between protected attributes, as they may reveal hidden sources of bias not apparent when attributes are examined individually. Figure 5.1(d) shows the average recidivism rates across subgroups defined by the combination of gender, race, and age. For the purpose of this exploratory analysis, age is categorised into three groups based on the 33rd and 67th percentiles of the overall age distribution: 0–27, 28–37, and 38 or above. Notably, African-American males under 27 years of age exhibit the highest average recidivism rates. Without fairness constraints, the model may learn and reproduce this data pattern, leading to subgroup disparities in predictions; that is, higher predicted recidivism rates for some groups.
Figure 5.1: Empirical average recidivism rate by protected attributes
@@ 5.2.4 Results and analysis
In insurance pricing, it is essential for the claim frequency to be unbiased at the portfolio level, so that the predicted claim frequency does not overestimate or underestimate the total claim frequency on average (Lindholm and Wüthrich, 2024). However, it is likely that our model is biased, given that we used neural networks with early stopping and included a regularisation term (Wüthrich, 2020). In that case, we suggest applying a bias correction method, such as the simple post-processing method introduced by Lindholm and Wüthrich (2024), to ensure the balance property of our predicted claim frequencies, so that the sum of predicted claim frequencies is the same as the sum of the observed claim frequencies. In theory, the balance property guarantees an unbiased claim frequency for any future portfolio whose risk profile, that is, the distribution of rating factors, matches that of the training data.
@@ Acknowledgments
All results in this paper are reproducible using Python, with code and data available on Zenodo (https://zenodo.org/records/16193169) and GitHub (https://github.com/agi-lab/dCov-fairness). The code was run in an Azure ML notebook on a Standard_E16s_v3 virtual machine with 16 cores, which supports the full pipeline but was used specifically for the hyperparameter tuning phase. The Azure notebook is available on Zenodo, while a Quarto-adapted version was used for the final runs and is provided on GitHub. For the COMPAS dataset, hyperparameter tuning took approximately 4 hours. For the pg15training dataset, tuning was stopped after 24 hours, as continuing would have required considerably more time and computational resources. This early stopping is not expected to materially affect the results, as the large dataset size supports good generalisation performance. The remaining parts of the pipeline (training and evaluation using the selected hyperparameters) were run in Quarto locally, taking approximately 5 hours for COMPAS and 7 hours for pg15training.
@@ Appendix A Details on distance covariance
where $\langle\mathbf{a},\mathbf{b}\rangle=\sum_{i=1}^{m}a_{i}b_{i}$ is the dot product of two vectors $\mathbf{a},\mathbf{b}\in\mathbb{R}^{m}$. Let $\mathbf{Y}$ be another random vector of dimension $n$. We say $\mathbf{X}$ and $\mathbf{Y}$ are independent if and only if their characteristic function satisfies
@@
which follows directly from the definition of independence (Billingsley, 2017). This motivates the development of a measure of distance between $\phi_{\mathbf{X}}(t)\phi_{\mathbf{Y}}(s)$ and $\phi_{\mathbf{X},\mathbf{Y}}(t,s)$. Let $f(\cdot)$ be a function with domain $\Omega$ and $w(\cdot)$ be a nonnegative weight function. The weighted $L_{2}$ norm of $f$ is defined as
@@
Using (A.3), Székely et al. (2007) developed the distance covariance for two random vectors $\mathbf{X}$ and $\mathbf{Y}$, which is the weighted $L_{2}$ norm of the difference between $\phi_{\mathbf{X},\mathbf{Y}}(t,s)$ and $\phi_{\mathbf{X}}(t)\phi_{\mathbf{Y}}(s)$:
@@
The unbiased estimator of the distance covariance, computed from a sample of $(\mathbf{X},\mathbf{Y})$ of size $n$, denoted by $(\mathbf{x},\mathbf{y})$, is defined as (Székely and Rizzo, 2014):
@@
with $|\cdot|$ denoting the Euclidean norm. The $\mathcal{U}$-centred matrix is specifically designed so that the resulting estimator of distance covariance is unbiased under independence, meaning its expectation is zero when the variables are independent (Székely and Rizzo, 2023). Furthermore, following the proof of Proposition 1 in (Székely and Rizzo, 2014), we can expand Equation (A.6) into an equivalent form:
@@ B.1 Mutual independence
Let $\mathbf{X}_{1},\ldots,\mathbf{X}_{d}$ be random vectors with dimensions $p_{1},\ldots,p_{d}$, taking values in $\mathbb{R}^{p_{1}},\ldots,\mathbb{R}^{p_{d}}$. Note that this also covers one-hot–encoded categorical variables, which can be treated as a multinoulli distribution (Bishop, 2006; Murphy, 2012).
@@
$p(\mathbf{X}_{i}=(x_{i,1},x_{i,2},...,x_{i,K})|\boldsymbol{\mu}=(\mu_{1},...,\mu_{K}))=\prod^{K}_{k=1}\mu_{k}^{x_{i,k}}.$ (B.1)
@@ B.2 Joint distance covariance
$dCov^{2}(\mathbf{X}_{\tau_{1}},.,\mathbf{X}_{\tau_{k}})$ (B.4)
$\qquad=\int_{\mathbb{R}^{p_{0}}}\left|\mathbb{E}\left[\prod^{k}_{i=1}\left(\phi_{\mathbf{X}_{\tau_{i}}}(\mathbf{x}_{\tau_{i}})-\exp\left\{\sqrt{-1}\langle\mathbf{x}_{\tau_{i}},\mathbf{X}_{\tau_{i}}\rangle\right\}\right)\right]\right|^{2}dw,$
with $dw=\prod^{k}_{i=1}\left(c_{p_{i}}dx_{\tau_{i}}\right)$ and $\mathcal{I}_{k}=\left\{(\tau_{1},\ldots,\tau_{k})\in\{1,\ldots,d\}^{k}:\tau_{1}<\tau_{2}<\cdots<\tau_{d}\right\}$ being the set of index tuples of size $d$. Intuitively, it is the summation of the pairwise associations ($dCov^{2}(\cdot,\cdot)$) and all the higher-order associations ($dCov^{2}(\cdot,\cdot,\cdots,\cdot)$) between the random vectors. The joint distance covariance characterises mutual independence if and only if this quantity is 0. To estimate this quantity from an available sample $\mathbf{x}_{1},\ldots,\mathbf{x}_{d}$ of size $n$, where each $\mathbf{x}_{i}\in\mathbb{R}^{p_{i}}$ corresponds to $\mathbf{X}_{i}$, Chakraborty and Zhang (2019) defined the bias-corrected estimator of the joint distance covariance as:
@@
with $\mathcal{I}\subseteq\{1,\ldots,d\}$ and $|\mathcal{I}|\geq 3$. In the setting of model predictions and $d\geq 2$ protected attributes, as in the regularisation term $\psi$ used in Equations (5.1) and (5.2), we have:
@@
where the term $\sum_{1\leq i,j\leq d}\widetilde{dCov}^{2}(\mathbf{s}_{i},\mathbf{s}_{j})+\zeta(\hat{\mathbf{y}},\mathbf{s}_{1},...,\mathbf{s}_{d})$ captures the association within the protected attributes and the extra associations as illustrated in the right panel of Figure 3.1.
@@
For notation simplicity, denote ${\mathbf{s}}=(\mathbf{s}_{1},\ldots,\mathbf{s}_{d})$ the sample of the protected attributes of size $n$ and ${\mathbf{s}}_{k}=(\mathbf{s}_{k,1},\ldots,\mathbf{s}_{k,d})$ be the $k$-th sample of $\mathbf{s}$}. By Equation (A.8), we can estimate the distance covariance between random vector of predictions $\hat{\mathbf{Y}}$ and the vector of protected attributes $\mathbf{S}$ as follows:
$\widetilde{dCov}^{2}(\hat{\mathbf{y}},{\mathbf{s}})=\frac{1}{n(n-3)}\Biggl[\sum_{k=1}^{n}\sum_{l=1}^{n}|\hat{\mathbf{y}}_{k}-\hat{\mathbf{y}}_{l}||{\mathbf{s}}_{k}-{\mathbf{s}}_{l}|$ (C.1)
@@
$\qquad-\biggl.\frac{2}{n-2}\sum_{k=1}^{n}\biggl(\sum^{n}_{l=1}|\hat{\mathbf{y}}_{k}-\hat{\mathbf{y}}_{l}|\biggr)\biggl(\sum^{n}_{l=1}|{\mathbf{s}}_{k}-{\mathbf{s}}_{l}|\biggr)\Biggr].$
@@
Here, each $\widetilde{dCov}^{2}(\hat{\mathbf{y}},\mathbf{s}_{i})$ is given by the unbiased estimator in Equation (A.8). Thus, Equation (C.6) can be understood as decomposing the joint distance covariance into the sum of the pairwise distance covariances and the residual intersection term $\eta(\hat{\mathbf{y}},\mathbf{s}_{1},\ldots,\mathbf{s}_{d})$.
The term $\eta(\hat{\mathbf{y}},\mathbf{s}_{1},\ldots,\mathbf{s}_{d})$ captures the dependence that arises only when the protected features are considered together. It is the residual link between the prediction $\hat{\mathbf{Y}}$ and the joint vector $\mathbf{S}=(\mathbf{S}_{1},\ldots,\mathbf{S}_{d})$ after removing the sum of the marginal dependencies with each $\mathbf{S}_{i}$. If $\eta$ is large, the prediction is reacting to the intersections of the protected attributes. In that case, even if the prediction and each protected attribute have a small distance covariance, indicating empirical independence, the predictions and the protected attributes are empirically mutually dependent. If $\hat{\mathbf{Y}}$ is independent of the whole vector $\mathbf{S}$, then the terms on the RHS of Equation (C.6) should be close to zero; i.e., $\hat{\mathbf{Y}}$ is independent of each protected attribute, and $\hat{\mathbf{Y}}$ is also independent of every subset of the protected attributes, leading to mutual independence.
Appendix D Details on model training and tuning